5. In flower beetles, pygmy (py) is recessive to normal size (py*), and red color (r) is recessive to brown (r). A beetle heterozygous for these characteristics was test crossed to a beetle homozygous for pygmy and red. The progeny phenotypes from this testcross are shown below. py 180 ey 22 Py 19 2x r191 Total412 What is the frequency of recombination between these two genes? Carry out a series of chi-square tests to determine if these genes are assorting independentl,y What does your chi-square value and associated probability indicate about whether these genes are linked?
Answer #1

1) py+ r+= 180

py+ r= 22

py r+= 19

py r= 191

Total=412

Recombination frequency= number of recombinant progeny/ total number of progeny X 100

The recombinants are py+ r and py r+

Recombination frequency= 22 +19/412= 0.09951 X 100= 9.951

Total proportion of py alleles= (19+191)= 210

b)

Expected ratio for this dihybrid cross is 8:8:8:8= 1:1:1:1 ratio

Degree of freedom= No of phenotypic characters-1= 4-1=3

The genes are not linked.

Chi square value at p=0.05 for df=3 is 12.838.

The (O-E) values are not identical in all columns.

Hence, the genes are not linked, as law of independent assortment is applicable (due to probability being more than 12.838).